# A Minimal Subset Approach for Efficient and Scalable Loop Closure

### v1

@@ IV-A Quantitative and Qualitative Analysis
For the batch experiments in Table I, the following pipeline is used: keyframes are sampled for all methods, and a pose graph is constructed using GTSAM [6] with odometry provided by KISS-ICP [37]. Potential loop closure matches are identified using OT descriptors, with a similarity threshold of 0.8. Ground truth data are used to classify candidates as true or false positives, based on a 1-meter radius. True positives are passed to small_gicp [38] for relative transformation estimation, with matches being verified if the registration residual is below 0.3 meters. Verified loop closures are then added as edges to the pose graph, which is optimized using the Levenberg-Marquardt algorithm.
Figure 5: Trajectory comparisons. Comparison of the ground truth and KISS-ICP initial poses, and the sampled poses after the pose graph optimization for the proposed Minimal Subset Approach (MSA) and the entropy-based approach on the KAIST sequence of the MulRan dataset and the SanJoseDowntown sequence of the Apollo-SouthBay dataset.
@@
Table I summarizes the results for all datasets and methods, where bold values represent the best performing adaptive method and underlined values indicate the best performing constant interval. The metrics include the Absolute Trajectory Error (ATE) for both translation and rotation, presented as the percent improvement in the KISS-ICP trajectory after pose graph optimization with each method. Additionally, the False Positive Ratio (FPR) for descriptor matching is provided, as sampling affects place recognition performance [30]. The results show that the proposed MSA consistently achieves the best performance, while constant intervals show varying results across different algorithms. In Fig. 4, the data from Table I are compiled into box plots, showing the deviation of each method’s ATE performance compared to using all samples, along with variations in memory usage and total execution time of the whole pipeline. The proposed approach maintains overall improvements with minimal performance loss in both translation and rotation, while significantly reducing memory usage and providing reasonable execution times. Although other methods reduce memory overhead and processing time, their performance varies considerably, lacking robustness. Finally, Figs. 1 and 5 visually compares ground truth poses, raw KISS-ICP poses, and the corrected trajectories for MSA and the entropy-based method after optimization. For all illustrated trajectories, the proposed approach consistently achieves lower translation and rotation RPE than the other methods, while the segments containing loop closures remain very close to the ground truth, demonstrating the approach’s ability to retain crucial keyframes.
IV-B Large-Scale Evaluation
TABLE II: Results for the Large-Scale Evaluation
Metrics t. ATE rot. ATE FPR MEM. TIME
All Samples nan nan nan > 128GB nan
Const. 1m 98.5% 98.4% 12% 90.9GB 7340s
Const. 2m 98.1% 97.3% 17% 47.2GB 2264s
Const. 3m 97.9% 97.0% 28% 35.4GB 1515s
Entropy 95.5% 93.4% 15% 51.0GB 4945s
Spacious. 94.9% 97.5% 24% 38.3GB 1752s
MSA (Ours) 98.3% 94.1% 12% 41.2GB 1951s
@@ IV-B Large-Scale Evaluation
Figure 6: Trajectory comparisons. Comparison of the ground truth and KISS-ICP initial poses, and the sampled poses after the pose graph optimization for the proposed Minimal Subset Approach (MSA) and the entropy-based approach on the SunnyvaleBigloop sequence of the Apollo-SouthBay dataset where the color gradient denotes the RPE. A grid block corresponds to 0.5x0.5 km2
For the second part, we conduct a more realistic online experiment using the SunnyvaleBigloop sequence from Apollo-SouthBay, covering over 100 kilometers. Poses are processed sequentially, simulating real-time operation. Keyframes are sampled online using the previously mentioned methods, appended to the pose graph, and optimized at each step with iSAM2 [8]. Simultaneously, descriptors are queried against past keyframes to detect potential loop closures, which are verified with small_gicp and integrated into the graph. This online approach provides insights into how well each sampling method handles large-scale missions. Table II presents the translation and rotation ATE, FPR, total allocated memory, and processing time. Notably, when using all samples, the system runs out of memory approximately 70% into the sequence, so no metrics are provided. The remaining sampling methods reduce memory usage while offering excellent performance. This is likely because the mission revisits the same locations multiple times, allowing frequent loop closure edges that prevent significant accumulated drift. In Fig. 6, the optimized trajectory of the entropy-based method is compared to MSA, with the color gradient indicating the translational RPE, showcasing the better performance while maintaining lower memory allocation and false positive ratio. As a final step, Figs. 7 and 8 provide further insights into execution time and memory allocation. Fig. 7 shows memory usage over time, illustrating how the sampling methods scale more efficiently compared to retaining all samples or using a constant 1-meter interval. Fig. 8 presents the moving average of loop closure detection time (left) and incremental pose graph optimization time (right). The top figures show constant interval methods, while the bottom shows adaptive methods. The proposed approach scales similarly to other adaptive methods while offering more robust performance, as demonstrated earlier. The shaded areas represent the minimum and maximum deviations, with the moving average trend indicating the challenge of increasing processing time in large-scale missions. While query time exceeds 1 second per new sample, and the incremental pose graph optimization reaches up to 0.15s for the constant 1-meter sampling method, MSA retains query times of less than 0.5s and incremental optimization times of less than 25ms, maintaining the real-time performance of the system throughout the large-scale mission.
@@
Figure 7: Memory allocation. Memory against mission time for the different sampling methods. Retaining all samples exceeds system memory of 128GB, demonstrating the need for efficient and scalable algorithms.
@@
Figure 8: Processing times. The top row shows constant interval methods, while the bottom row shows adaptive methods. The left column represents loop closure detection time, and the right column represents incremental pose graph optimization convergence time. All times in seconds (s).
V Conclusions
@@
In conclusion, as the robotics community advances towards larger-scale and long-term missions, addressing system limitations with efficient and scalable solutions becomes crucial. This article introduced the Minimal Subset Approach (MSA) to address the combined challenges of loop closure detection and pose graph optimization by reducing redundant samples while maintaining robust performance. Experimental evaluations across multiple datasets show that MSA outperforms in terms of FPR, as well as achieving superior ATE and RPE after pose graph optimization. Unlike other adaptive sampling methods, MSA requires minimal to no parameter tuning, while using a fraction of the system’s memory and maintaining comparable computational times.

### v2

@@ IV-A Quantitative and Qualitative Analysis
For the batch experiments in Table I, we evaluate our approach on three datasets: KITTI Odometry [31], MulRan [32], and Apollo-SouthBay [33], representing diverse environments, including urban areas, rural landscapes, and complex structures like bridges and tunnels. Keyframes are sampled for all methods, and the pose graph is constructed using GTSAM [7] with odometry provided from KISS-ICP [34]. Loop closure candidates are identified using OT or SC descriptors with a similarity threshold of 0.8. Candidates within a 3-meter radius of the ground truth are classified as true positives and processed with small_gicp [35] to estimate relative transformations. Matches with registration residuals below 0.3 meters are verified and added as edges to the pose graph, which is optimized using the Levenberg-Marquardt algorithm.
@@
Table I summarizes the results for all datasets, descriptors, and methods. The metrics include the Absolute Trajectory Error (ATE) for both translation and rotation, presented as the percent improvement in the KISS-ICP trajectory after pose graph optimization with each method. Additionally, the False Positive Ratio (FPR) for descriptor matching is reported, as sampling impacts place recognition performance [6]. The results demonstrate that the proposed MSA consistently delivers the best performance, while constant intervals yield varying results across different environments, highlighting the absence of a universal fixed sampling interval. In Fig. 3, the data from Table I are presented as box plots, illustrating the deviation of each method’s ATE performance compared to using all samples, alongside variations in memory usage and total pipeline execution time. The proposed approach sustains overall improvements with minimal performance loss in both translation and rotation, while significantly reducing memory usage and execution time. Although other methods also reduce memory and processing overhead, their performance lacks robustness and varies considerably. Finally, Figs. 1 and 4 compare ground truth poses, raw KISS-ICP poses, and corrected trajectories for MSA and the entropy-based method post optimization. Across all illustrated trajectories, the proposed approach achieves consistently lower translation and rotation RPE than other methods, with loop closure segments closely aligning with the ground truth, demonstrating its ability to retain essential edges.
@@
Figure 4: Trajectory comparisons. Comparison of the ground truth, raw KISS-ICP, and the sampled poses after the pose graph optimization for the proposed Minimal Subset Approach (MSA) and the entropy-based approach on the KAIST sequence of the MulRan dataset and the SanJoseDowntown sequence of the Apollo-SouthBay dataset.
IV-B Larger-Scale Evaluations
TABLE II: Results for the Larger-Scale Evaluations and the Ablation Study for the proposed approach
Metrics t. ATE R. ATE FPR MEM. TIME
Apollo-SB Sunnyvale All Samples nan nan nan >128GB nan
Constant 1.0m 98.5% 98.4% 12% 90.9GB 7340s
Constant 2.0m 98.1% 97.3% 17% 47.2GB 2264s
Constant 3.0m 97.9% 97.0% 28% 35.4GB 1515s
Entropy-based 95.5% 93.4% 15% 51.0GB 4945s
Spaciousness 94.9% 97.5% 24% 38.3GB 1752s
MSA (default) 98.3% 94.1% 12% 41.2GB 1951s
MSA (w/o $\pi_{\tau}$) 94.7% 93.2% 13% 32.4GB 1438s
MSA (w/o $\rho_{\tau}$) 98.9% 96.4% 12% 43.7GB 2022s
MSA ($\alpha=10$) 96.5% 93.9% 12% 37.6GB 1764s
MSA ($\beta=10$) 98.6% 95.8% 12% 42.0GB 2004s
Newer-College Park All Samples 93.9% 78.9% 02% 65.8GB 6742s
Constant 0.1m 89.7% 71.4% 03% 63.2GB 4987s
Constant 0.3m 79.9% 58.5% 08% 28.9GB 1262s
Constant 1.0m 60.7% 43.1% 23% 09.2GB 165s
Entropy-based 87.3% 65.8% 05% 26.3GB 1401s
Spaciousness 72.5% 52.5% 14% 27.2GB 1096s
MSA (default) 90.1% 67.0% 02% 38.7GB 1884s
MSA (w/o $\pi_{\tau}$) 81.7% 41.4% 03% 12.8GB 346s
MSA (w/o $\rho_{\tau}$) 97.7% 81.4% 02% 64.7GB 6226s
MSA ($\alpha=10$) 86.9% 64.4% 02% 29.2GB 1498s
MSA ($\beta=10$) 93.2% 69.4% 02% 42.5GB 2791s
To support our last contribution, we conduct more realistic, larger-scale, online experiments. Poses are processed sequentially, simulating real-time operation. Keyframes are sampled online using the previously mentioned methods, appended to the pose graph, and optimized at each step with iSAM2 [8]. Simultaneously, OT descriptors are queried against past keyframes to detect loop closures, which are verified with small_gicp and integrated into the graph.
IV-B1 Urban Driving Scenario
For this scenario we use one of the datasets from the batch experiments, Apollo-SouthBay, focusing on the SunnyvaleBig- loop sequence, which spans over 100 kilometers. Table II reports the translation and rotation ATE, FPR, total allocated memory, and processing time. Notably, when using all samples, the system exhausts memory, approximately 70% into the sequence, preventing the calculation of the metrics. The other sampling methods reduce memory usage while maintaining excellent performance. This is likely due to the sequence revisiting the same locations multiple times, enabling frequent loop closure edges that mitigate accumulated drift. In Fig. 5, the optimized trajectory of the entropy-based method is compared to MSA, with a color gradient representing translational RPE. The visualization highlights MSA’s superior performance, achieved with lower memory allocation and a reduced false positive rate. Lastly, Fig. 6 shows the memory usage over time, illustrating how sampling methods scale more efficiently compared to retaining all samples or using a constant 1-meter interval.
Apollo-SouthBay Sunnyvale
Newer-College Park
Figure 5: Trajectory comparisons. The ground truth and raw KISS-ICP poses compared against the sampled and optimized poses using the proposed Minimal Subset Approach (MSA) and the entropy-based method for the Sunnyvale and the constant 1-meter approach for the Park. The color gradient represents the Relative Pose Error (RPE). Grid spacing corresponds to 0.5x0.5 km2 for the left figure and 50x50 m2 for the right figure.
IV-B2 Campus Scenario
To provide additional insights, we evaluate a different dataset, the Newer-College [36], and more specifically the Park sequence, a 30-minute mapping session with multiple loops. This scenario, resembling a mobile ground robot, differs significantly from Apollo-SouthBay due to its more confined environment and denser frame sampling, approximately 0.1 meters compared to 1 meter. These differences highlight how distinct missions impose unique requirements.
From Table II, constant intervals show that, a 1-meter sampling, severely underperforms in this scenario, retaining only 7% of the frames. This leads to significant performance losses in both ATE and FPR. To address this, we include lower constant intervals of 0.1 and 0.3 meters for comparison. While these improve performance, our proposed MSA outperforms them by prioritizing informative sampling, providing more valuable nodes to the pose graph. The entropy-based method closely follows, emphasizing the importance of spatial features in sampling. Here, we note that both the spaciousness- and entropy-based methods required additional tuning for this scenario. As in the previous experiment, Fig. 5 compares the optimized trajectory of MSA with the 1-meter approach, using a color gradient to depict translational RPE, which highlights MSA’s superior metric localization performance. Figs 6 and 7 further compare memory allocation over time and processing times. Fig. 7 presents box plots for incremental PGO (left) and Loop Closure Detection (right) processing times per step. Notably, both tasks exhibit significantly high processing times, potentially causing delays in real-world missions. While lower fixed intervals (0.1 and 0.3 meters) reduce processing time, 0.3-meter and 1-meter intervals show subpar performance, and the 0.1-meter interval demands substantial memory allocation since it is very close to the native sampling of the dataset. Interestingly, methods achieving good ATE performance, such as the entropy-based and MSA, demonstrate higher outliers in the incremental PGO, suggesting that valuable edges are added to the pose graph, contributing to their better performance.
@@ IV-B2 Campus Scenario
Figure 6: Memory allocation. Memory against each acquired frame for the different sampling methods. Retaining all samples exceeds system memory of 128GB, demonstrating the need for efficient and scalable algorithms.
@@
Figure 7: Processing times. The distributions of the incremental pose graph optimization convergence (left), and the loop closure detection (right) time, per new frame, for the Park sequence of the Newer-College dataset.
IV-C Ablation Study
In all experiments, we maintained constant hyperparameters ($\alpha=\beta=1$), demonstrating that the proposed method adapts seamlessly to different environments and descriptors without requiring manual tuning. To further analyze the impact of each optimization term, we conducted an ablation study, with results shown in Table II. Specifically, we tested scenarios where only one term, $\rho_{\tau}$ or $\pi_{\tau}$, was used, and scenarios where the weight of one term was increased to 10 while keeping the other at 1. The results indicate that emphasizing the information preservation term or using it exclusively improves performance but leading to increased number of keyframes sampled. Conversely, prioritizing redundancy minimization significantly reduces memory usage but negatively impacts performance. Based on these, practitioners can decide whether to prioritize memory or performance, depending on their specific application requirements.
V Discussion
The results highlight that no universal constant sampling interval is suitable for all scenarios. While a constant 1-meter interval performs well in datasets similar to the Urban Driving Scenario, showing minimal deviations in Fig. 3, it performed poorly in the mobile robotics Campus Scenario. This reinforces our core argument: adaptive methods that account for both place recognition performance and pose graph optimization are essential. Naive methods that rely on spatial heuristics, such as spaciousness- and entropy-based approaches, require extensive tuning and prior environmental knowledge, limiting their applicability. Our proposed MSA consistently excels across all scenarios without the need for parameter tuning, demonstrating robustness in this traditional loop closure setup. It is effective with both widely used handcrafted and learning-based descriptors, showcasing adaptability to different datasets and descriptors. As future work, we aim to test our method with full global localization pipelines, such as RING++ [37] or LCDNet [38].
VI Conclusions
@@
In conclusion, as the robotics community advances towards larger-scale and long-term missions, addressing system limitations with efficient and scalable solutions becomes crucial. This article introduced the Minimal Subset Approach (MSA) to address the combined challenges of loop closure detection and pose graph optimization by reducing redundant samples while maintaining robust performance. Experimental evaluations across multiple datasets show that MSA excels in terms of FPR, as well as in achieving superior ATE and RPE after pose graph optimization. Unlike other adaptive sampling methods, MSA requires minimal to no parameter tuning, while using a fraction of the system’s memory and maintaining comparable computational times.